Assume sales for Peach Street Industries are expected to increase by 8.00% from 2015 to 2016. Peach Street is operating at full capacity currently and expected assets-to-sales and spontaneous liabilities-to-sales to remain the same. Additionally, the firm is looking to maintain their 2015 net profit margin and dividend payout ratios for 2016. The firm’s tax rate is 37.00% and selected income statement and balance sheet information for 2015 is provided below:

Entry Value Entry Value
Current Assets $800.00 Sales $2,500.00
Net Fixed Assets (NFA) $700.00 Operating Costs $2,030.00
Total Assets $1,500.00 Depreciation $90.00
Accounts Payable and Accruals $30.00 Interest Expense $69.00
Notes Payable $180.00 Dividends Paid $93.30
Long term debt $510.00
Total Equity $780.00

How much in additional funds (external capital) will Peach Street Industries need in 2016 to support their projected growth in sales? (i.e., calculate the firm’s additional funds needed --AFN)
Answer #1

Additional Funds Needed = [A0 x (ΔS / S0)] - [L0 x (ΔS / S0)] - [S1 x PM x b]

Where,
Ao = current level of assets
Lo = current level of liabilities
ΔS/So = percentage increase in sales i.e. change in sales divided by current sales
S1 = new level of sales
PM = profit margin
b = retention rate = 1 - payout rate

Net Income = [Sales - Operating Costs - Depreciation - Interest Expense] * [1 - t]

= [2,500 - 2,030 - 90 - 69] * [1 - 0.37] = 311 * 0.63 = $195.93

Profit Margin = Net Income / Sales = $195.93 / $2,500 = 0.078372, or 7.8372%

Retention Ratio = 1 - Payout Rate = 1 - (Dividend Paid / Net Income)

= 1 - (93.30 / 195.93) = 1 - 0.4762 = 0.5238, or 52.38%

AFN = [$1,500 x 0.08] - [$30 x 0.08] - [{$2,500 * (1 + 0.08)} x 0.078372 x 0.5238]

= $120 - $2.40 - $110.84 = $6.7596